OpenStudy (anonymous):

write each equation in standard form : A>0 1/2 (x+1)+1/5 (y-2) +1=0

OpenStudy (anonymous):

\[.5x+.5+\frac{ 1 }{ 5 }y-\frac{ 2 }{ 5 }+1=0\]

OpenStudy (anonymous):

i think your answer is wrong...?

OpenStudy (anonymous):

then \[.5x+\frac{ 1 }{ 5 }y+(.5+1-\frac{ 2 }{ 5 })=0\]

OpenStudy (anonymous):

\[.5x+\frac{ 1 }{ 5 }y+1.1=0\]

OpenStudy (anonymous):

can u explain your answer .?

OpenStudy (anonymous):

which standard equation you want ? y=mx+c ??

OpenStudy (anonymous):

ah.yup,im here. just wait.

OpenStudy (anonymous):

try this one .you need to write each equation in slope intercept form. x/3=1-y/2

OpenStudy (anonymous):

and this one ..x+2y-1=3x-1/2

OpenStudy (anonymous):

the slope intercept form is y=mx+c the first one will be \[y=\frac{ -2 }{ 3 }x+2\] where slope=-2/3 y-intercept=2

OpenStudy (anonymous):

te second one will be \[y=x+\frac{ 1 }{ 4 }\] where slope=1 y-intercept=1/4

OpenStudy (anonymous):

Ax+By+C=0.equation form ..1/2 (x+1)+1/5 (y-2) +1=0 can u try this one?/

OpenStudy (anonymous):

first distribute 1/2 and 1/5

OpenStudy (anonymous):

how? i dont get it.

OpenStudy (anonymous):

\[.5x+.5+\frac{ 1 }{ 5 }y-\frac{ 2 }{ 5 }+1=0\]

OpenStudy (anonymous):

ammmh.okei . thank you .!! very much..

OpenStudy (anonymous):

can you complete the answering ?

OpenStudy (anonymous):

is'nt complete?

OpenStudy (anonymous):

you have to collect the constants .5,1,-2/5